OpenStudy (anonymous):

Write a point-slope equation for the line that passes through the point (1, 6) and is parallel to the line given by y = 5x + 25. Do not use parenthesis on the y-side

OpenStudy (anonymous):

The attachment should help :)

OpenStudy (anonymous):

thks

OpenStudy (anonymous):

tell me if u have any questions when u solve it :)

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

i think i have but i cant get it wrong y=6x+1 is what i came up with

OpenStudy (anonymous):

k let me check

OpenStudy (anonymous):

Close :) Ok so u got the wrong slope but the rest was right, think u know what u did wrong?

OpenStudy (anonymous):

would the slope be 5? and im not sure what i did wrong

OpenStudy (anonymous):

yeah the slope is 5, i think it might have been a careless mistake that u accidentally put 6 instead of 5 as the slope, don't worry I do that all the time :)

OpenStudy (anonymous):

okay heres a new question (-1, 1) and (-3,-5) A. y + 3 = 3(x + 5) B. y = 3x + 4 C. y - 1 = 3(x + 1) D. y + 5 = 3(x + 3)

OpenStudy (anonymous):

it gives u this with no slope? just making sure

OpenStudy (anonymous):

nvm seems like u posted it as a new question & someone's already answering it

OpenStudy (anonymous):

the slope would be 3

OpenStudy (anonymous):

got the question right but failed my work sheet

OpenStudy (anonymous):

oh ok, need help with the others?

OpenStudy (anonymous):

ya if u wanna help me

OpenStudy (anonymous):

sure post it as a new question it'll be easier :)